# MOGO: Residual Quantized Hierarchical Causal Transformer for High-Quality and Real-Time 3D Human Motion Generation

### v1

@@ Abstract
Recent advances in transformer-based text-to-motion generation have led to impressive progress in synthesizing high-quality human motion. Nevertheless, jointly achieving high fidelity, streaming capability, real-time responsiveness, and scalability remains a fundamental challenge. In this paper, we propose MOGO (Motion Generation with One-pass), a novel autoregressive framework tailored for efficient and real-time 3D motion generation. MOGO comprises two key components: (1) MoSA-VQ, a motion scale-adaptive residual vector quantization module that hierarchically discretizes motion sequences with learnable scaling to produce compact yet expressive representations; and (2) RQHC-Transformer, a residual quantized hierarchical causal transformer that generates multi-layer motion tokens in a single forward pass, significantly reducing inference latency. To enhance semantic fidelity, we further introduce a text condition alignment mechanism that improves motion decoding under textual control. Extensive experiments on benchmark datasets including HumanML3D, KIT-ML, and CMP demonstrate that MOGO achieves competitive or superior generation quality compared to state-of-the-art transformer-based methods, while offering substantial improvements in real-time performance, streaming generation, and generalization under zero-shot settings.
1 Introduction
@@
Despite these achievements, both diffusion- and LLM-based frameworks face practical limitations [7, 8, 9, 10]. Diffusion models often rely on iterative refinement processes, which introduce significant inference latency and hinder their suitability for real-time or interactive applications [7]. On the other hand, LLM-based models, although autoregressive, typically involve large parameter sizes and long context dependencies, leading to high memory and computation costs that challenge deployment on lightweight or streaming scenarios [9]. A broader discussion of related research trends is provided in Section 4.
To address these issues, we propose MOGO (Motion Generation with One-pass), a transformer-based autoregressive framework that directly predicts motion tokens in a single forward pass. Compared to diffusion-based models, MOGO enables streamable motion decoding without iterative sampling, significantly reducing inference time. Furthermore, unlike LLM-based motion generators, MOGO adopts a lightweight yet expressive transformer architecture with considerably fewer parameters—allowing for real-time token generation while maintaining high motion fidelity.
MOGO is composed of two key components: (1) MoSA-VQ (Motion Scale-Adaptive Residual VQ-VAE), which discretizes motion sequences via hierarchical quantization and learnable scaling, producing compact and adaptive representations [11]; and (2) RQHC-Transformer (Residual Quantized Hierarchical Causal Transformer), which autoregressively generates multi-layer motion tokens in a single pass [12]. As shown in Figure 1, this design allows for the synthesis of coherent and extended motion sequences while preserving streaming capability and generalization across diverse prompts and datasets.
Extensive experiments demonstrate that MOGO achieves superior results on both in-distribution and out-of-distribution scenarios, outperforming state-of-the-art LLM-based motion generators in quality, inference speed, and scalability—making it a robust and real-time-capable solution for practical text-to-motion generation.
Our main contributions:
MOGO: A unified transformer-based framework for one-pass high-fidelity motion generation. It achieves generation quality comparable to state-of-the-art LLM- and diffusion-based methods, while significantly improving real-time efficiency and enabling streamable motion synthesis.
MoSA-VQ: A motion discretization module that employs hierarchical quantization and learnable feature scaling for compact and expressive representations.
RQHC-Transformer: A lightweight hierarchical causal transformer that autoregressively generates multi-layer motion tokens in one forward pass.
All code and demonstration pages will be made publicly available on our GitHub repository upon acceptance.
2 Methods
We introduce MOGO, a unified framework designed for efficient and expressive text-to-motion generation. Mogo could generate a high-quality motion sequence from a textual description, where each frame represents a -dimensional human pose. As shown in Figure 2, our MOGO framework consists of two core components: (1) MoSA-VQ (left) that discretizes motion into multi-level tokens with high precision, and (2) RQHC-Transformer (right) leveraging hierarchical causal forward computation to generate motion tokens collectively in a streamlined manner (algorithms are shown in Appendix C). To enhance zero/few-shot performance, we incorporate text condition alignment (TCA) for improved motion decoding.
2.1 MoSA-VQ: Motion Scale-Adaptive Residual VQ-VAE
To effectively transform continuous human motion sequences into discrete representations suitable for autoregressive modeling, we adopt a residual vector quantized variational autoencoder (RVQ-VAE) architecture. This module, termed MoSA-VQ, serves as the foundation for learning compact yet expressive motion codes. Residual quantization enables progressive refinement of motion features across multiple levels, capturing both coarse and fine-grained temporal dynamics essential for high-fidelity generation.
However, previous RVQ-based approaches [2, 5] often rely on manually calibrated feature ranges, which limit their adaptability and quantization efficiency in complex motion settings. To address this, we introduce a novel learnable feature scaling mechanism, allowing the quantization space to dynamically adapt at each hierarchy level. This design not only eliminates the need for manual tuning but also improves reconstruction accuracy and codebook utilization.
Hierarchical Quantization with Feature Scaling.
Motion encoder outputs a latent representation , where is the number of motion tokens and is the feature dimension. At each level , we perform feature scaling before residual quantization. Let be the residual feature at level 0. We first compute a scaled version of the residual feature:
where and are learnable scale and bias parameters. The input is the residual feature passed from the previous quantization level and represents the information that has not yet been captured.
We then apply vector quantization to the scaled residual:
where maps each input to the nearest codebook entry. This quantized output exists in the scaled feature space, so we invert the transformation to project it back into the original residual space:
The recovered quantized vector captures the portion of the motion representation explained at level , and we update the residual for the next layer by subtracting this contribution:
This process repeats hierarchically across all levels, resulting in a set of quantized vectors . The final quantized latent representation is reconstructed by summing contributions from all layers:
which is subsequently passed to the decoder to generate the full motion sequence.
@@ 2.1 MoSA-VQ: Motion Scale-Adaptive Residual VQ-VAE
To guide the hierarchical quantization process toward informative and stable motion representations, we jointly optimize a multi-term objective designed to promote accurate reconstruction, effective codebook usage, and numerical stability. The overall loss function is defined as:
where denotes the stop-gradient operator. The first term encourages accurate motion reconstruction, the second enforces codebook commitment, and the third stabilizes training by constraining the scale and bias parameters to remain close to identity mappings.
To preserve fine-grained temporal details throughout the encoding and decoding process, we employ stride-1 convolutions in both the encoder and decoder. This design avoids excessive downsampling, which can introduce high-frequency aliasing artifacts and degrade motion continuity.
Learnable scaling allows the quantization layers to focus on encoding residuals at different abstraction levels, with deeper layers specializing in finer motion refinements. Such a design increases the expressiveness and interpretability of the latent representation, contributing to higher reconstruction quality and better hierarchical decomposition of motion dynamics.
2.2 RQHC-Transformer: Residual Quantized Hierarchical Causal Transformer
multi-level token representation via MoSA-VQ, the next challenge is to generate these tokens from a textual prompt in a temporally coherent and semantically aligned manner. To this end, we propose the RQHC-Transformer, a specialized autoregressive model designed to synthesize motion token sequences across multiple quantization levels in a single unified framework.
Unlike prior transformer-based models that process motion tokens at a single resolution or require separate modules per layer [2], our RQHC-Transformer embraces the hierarchical structure of the quantized latent space. It sequentially predicts tokens from coarse to fine levels, enabling the model to first capture the overall motion structure and then refine it with increasing detail. This mirrors the residual nature of MoSA-VQ and preserves the interpretability of each quantization stage.
Input Construction.
@@ 2.2 RQHC-Transformer: Residual Quantized Hierarchical Causal Transformer
where and are the query and key embeddings at position and , and is a learned relative position embedding, and and are learnable global bias vectors.
This formulation allows the model to incorporate relative distance information into the attention computation. The final output of the masked self-attention is:
@@
Autoregressive Objective.
We use a multi-layer conditional maximum likelihood objective:
Here, is the batch size, is the sequence length for the -th sample, and is the number of quantization levels. denotes the ground truth discrete token at position in layer , while represents the model’s predicted context (e.g., previously generated tokens or hidden states) up to position for sample and layer . The conditional input encodes auxiliary information such as the text prompt, and denotes the model’s predicted probability distribution parameterized by weights .
2.3 Text Condition Alignment for Improved Motion Decoding
Despite the RQHC-Transformer being trained on structured datasets such as HumanML3D, real-world user prompts at inference time often exhibit substantial variation in phrasing, granularity, and semantic clarity. This distribution shift between training and deployment introduces a critical challenge: the model may underperform on out-of-distribution inputs due to linguistic mismatch between test-time instructions and training-time prompts.
To mitigate this discrepancy, we propose TCA (text condition alignment), a lightweight inference-time module that reformulates raw user instructions into a training-compatible form without altering the model’s weights. TCA operates by leveraging a pretrained language model , which serves as a prompt rewriting transformer to align free-form inputs with the motion model’s learned linguistic prior.
We denote the space of unstructured user prompts as , which includes arbitrary, potentially noisy natural language descriptions. In contrast, denotes the domain of structured, semantically aligned prompts used during training. Given a test-time input , TCA produces an aligned instruction by querying the rewriting model under a style prior :
where encapsulates syntactic and semantic formatting patterns derived from the training corpus.
The aligned prompt is then directly used as the conditioning input for the motion generation pipeline, replacing the original . This inference-time alignment improves compatibility between the textual condition and the model’s learned text-motion mapping, enabling more robust, coherent, and semantically accurate generation—particularly in zero- and few-shot settings. Since TCA operates independently of model parameters and requires no additional supervision, it offers a practical and efficient solution for open-domain deployment.
3 Experiments
3.1 Datasets and Evaluation Metrics
Datasets. We train and evaluate MOGO on HumanML3D [13] and KIT-ML [14]. HumanML3D contains 14,616 motion clips and 44,970 textual annotations, while KIT-ML includes 3,911 motion clips and 6,278 descriptions. Data splits follow T2M [13] with ratios of 0.8/0.15/0.05. To assess out-of-distribution generalization, we evaluate zero-shot performance on the CMP dataset [15], which includes 8,700 non-daily, combat-style motions. CMP provides three levels of text descriptions, and we use its test split. During CMP evaluation, prompt engineering is disabled to ensure fairness.
Evaluation Metrics. We adopt standard metrics from [13]. FID (Frechet Inception Distance) measures motion realism and serves as a primary evaluation metric. R-Precision evaluates text-motion alignment (Top-1, 2, 3). MM-Dist computes the distance between motion and text embeddings. MultiModality measures the variance of motions generated from the same text prompt.
3.2 Real-time Inference
To enhance zero- and few-shot generalization, we leverage GLM-4 [16] to rewrite user prompts that deviate from the HumanML3D style, aligning them more closely with the model’s learned text-motion priors. Moreover, benefiting from MOGO’s autoregressive design, motion tokens are generated in a streaming fashion—each token is predicted and emitted sequentially—enabling low-latency, frame-by-frame output without the need to wait for the entire sequence.
This architectural advantage allows MOGO to operate in real-time settings, achieving an inference speed of 30 frames per second on a single NVIDIA A100 GPU, making it well-suited for interactive applications such as virtual avatars, animation systems, and embodied agents.
3.3 Comparison to State-of-the-art Approaches
3.3.1 Reconstruction Quality of Motion Encoder
We compare the reconstruction fidelity of our MoSA-VQ with other state-of-the-art motion VAEs. As shown in Table 1, our encoder achieves the lowest FID on both HumanML3D and KIT-ML datasets.
| HumanML3D | KIT-ML | ||
| Method | FID | Method | FID |
| M2DM [17] | 0.063±0.001 | M2DM [17] | 0.413±0.009 |
| T2M-GPT [3] | 0.070±0.001 | T2M-GPT [3] | 0.472±0.011 |
| MoMask [2] | 0.019±0.001 | MoMask [2] | 0.112±0.002 |
| MMM [4] | 0.075±0.001 | MMM [4] | 0.641±0.014 |
| MOGO | 0.013±0.001 | MOGO | 0.037±0.001 |
3.3.2 Quantitative Results
As shown in Table 2, MOGO exhibits consistently strong performance across the HumanML3D, KIT-ML, and CMP (zero-shot) benchmarks. With the incorporation of our proposed TCA module, MOGO achieves leading results in several key metrics—particularly in terms of generation quality and diversity—while maintaining competitive performance across all others.
On in-distribution datasets like HumanML3D and KIT-ML, MOGO with TCA demonstrates top-tier performance across most metrics, though it exhibits a slight increase in FID compared to the base MOGO, the overall generation quality remains state-of-the-art, underscoring the robustness of our hierarchical architecture.

### v2

# MOGO: Residual Quantized Hierarchical Causal Transformer for Real-Time and Infinite-Length 3D Human Motion Generation

@@ Abstract
Recent advances in transformer-based text-to-motion generation have significantly improved motion quality. However, achieving both real-time performance and long-horizon scalability remains an open challenge. In this paper, we present MOGO (Motion Generation with One-pass), a novel autoregressive framework for efficient and scalable 3D human motion generation. MOGO consists of two key components. First, we introduce MoSA-VQ, a motion scale-adaptive residual vector quantization module that hierarchically discretizes motion sequences through learnable scaling parameters, enabling dynamic allocation of representation capacity and producing compact yet expressive multi-level representations. Second, we design the RQHC-Transformer, a residual quantized hierarchical causal transformer that decodes motion tokens in a single forward pass. Each transformer block aligns with one quantization level, allowing hierarchical abstraction and temporally coherent generation with strong semantic flow. Compared to diffusion- and LLM-based approaches, MOGO achieves lower inference latency while preserving high motion fidelity. Moreover, its hierarchical latent design enables seamless and controllable infinite-length motion generation, with stable transitions and the ability to adaptively incorporate updated control signals at arbitrary points in time. To further enhance generalization and interpretability, we introduce Textual Condition Alignment (TCA), which leverages large language models with Chain-of-Thought reasoning to bridge the gap between real-world prompts and training data. TCA not only improves zero-shot performance on unseen datasets but also enriches motion comprehension for in-distribution prompts through explicit intent decomposition. Extensive experiments on HumanML3D, KIT-ML, and the unseen CMP dataset demonstrate that MOGO outperforms prior methods in generation quality, inference efficiency, and temporal scalability.
I Introduction
@@
Despite these achievements, both diffusion- and LLM-based frameworks face practical limitations [7, 8, 9, 10]. Diffusion models often rely on iterative refinement processes, which introduce significant inference latency and hinder their suitability for real-time or interactive applications [7]. On the other hand, LLM-based models, although autoregressive, typically involve large parameter sizes and long context dependencies, leading to high memory and computation costs that challenge deployment on lightweight scenarios [9]. A broader discussion of related research trends is provided in Section II.
To address the challenges of motion generation quality, inference efficiency, and generalization, we propose MOGO, a unified and efficient transformer-based framework for expressive text-to-motion synthesis. MOGO generates high-fidelity motion sequences from textual input in a single forward pass, combining structural compactness, temporal coherence, and extendable decoding. The framework is composed of four key components, each targeting a core aspect of the generation pipeline:
We propose MoSA-VQ, a motion structure-aware quantization framework with two key innovations: (i) a learnable modulation mechanism that adaptively scales residual magnitudes across quantization levels, and (ii) a novel cross-level decorrelation loss that enforces statistical orthogonality between quantized vectors and subsequent residuals. Together, these components enable more compact and hierarchically disentangled representations, leading to better reconstruction under constrained codebook capacity.
RQHC-Transformer (Residual Quantized Hierarchical Causal Transformer) functions as the motion decoder. It autoregressively generates multi-layer motion tokens in a single forward pass, with each transformer block dedicated to decoding one level of the quantized hierarchy. This design ensures temporally coherent synthesis while preserving decoding efficiency and compatibility with interactive applications.
Infinite-Length Generation is enabled by the autoregressive nature of MOGO’s architecture. Unlike diffusion-based methods that require fixed-length sampling, MOGO supports seamless motion extension with consistent structure and fidelity, making it suitable for real-time or open-ended generation scenarios.
Textual Condition Alignment (TCA) leverages large language models with Chain-of-Thought reasoning to bridge the gap between real-world user prompts and the model’s training distribution. Beyond simple style normalization, TCA explicitly decomposes complex motion instructions into structured semantic cues, enabling more accurate motion interpretation. This enhances MOGO’s generalization capability across both in-distribution and zero-/few-shot scenarios, while improving control fidelity and robustness to diverse user inputs.
II Related Work
LLM-Based Motion Generation Models. LLM-based architectures have become a cornerstone of text-to-motion generation, leveraging their ability to model sequential data and adapt to varied tasks [2, 3, 4, 5, 22]. Models like MoMask [2] and MMM [4] use masked token modeling to produce high-quality motions but struggle with extendable output and generalization in low-data or out-of-distribution settings due to their bidirectional design. Conversely, autoregressive models like T2M-GPT [3] and AttT2M [5] enable sequential generation, making them suitable for real-time applications and scalable with larger datasets, though they often sacrifice some motion quality. Efforts like MotionGPT [22] integrate multimodal language modeling but face challenges in achieving high-fidelity motion outputs. Our MOGO framework addresses these issues by combining efficient encoding through the MoSA-VQ with a single-pass autoregressive transformer, ensuring both high-quality motion and interactive capabilities.
III Methods
III-A Overview of the MOGO Framework
Our proposed framework, MOGO, as showned in Figure 1, consists of four key components designed for efficient and expressive text-to-motion generation. First, MoSA-VQ discretizes motion sequences into multi-level latent tokens using hierarchical residual quantization with adaptive scaling. Second, RQHC-Transformer autoregressively decodes the quantized representations with a stack of level-aligned transformer blocks. Third, MOGO supports infinite-length generation via seamless autoregressive extension. Lastly, we introduce TCA to enhance generalization to real-world prompts. We elaborate on each component in the following subsections.
III-B MoSA-VQ: Motion Scale-Adaptive Residual VQ-VAE
To transform continuous human motion into discrete representations suitable for autoregressive generation, we build upon residual vector quantized variational autoencoders (RVQ-VAEs) and propose MoSA-VQ. While standard RVQ-based approaches [2, 5] hierarchically encode motion features through fixed-stage quantization, they often rely on manually tuned feature ranges and static scaling behavior, which limit their adaptability and quantization efficiency—especially in the presence of diverse and complex motions.
Adaptive Hierarchical Quantization.
To address these limitations, we introduce a learnable feature scaling mechanism into each quantization level. Specifically, each level is equipped with its own learnable scale and bias parameters, which dynamically adjust the amplitude and offset of the input residual features before quantization. This adaptive design allows each level to allocate representational capacity where needed, resulting in more effective codebook usage and higher reconstruction accuracy without the need for handcrafted calibration.
At each hierarchical quantization level , the residual vector is first transformed using a learnable affine mapping. We denote this process as:
where is a diagonal scaling matrix, is a learnable bias vector, denotes nearest-neighbor quantization via a codebook, and represents the decoded contribution at level .
This hierarchical quantization scheme enables progressive residual refinement. Unlike conventional residual vector quantization (RVQ) frameworks that apply uniform scaling across all levels, our design introduces level-specific adaptive modulation through learnable parameters and . This flexibility allows coarse levels to capture global structure, while deeper layers focus on subtle motion or spatial details. The inverse affine transformation ensures consistent feature recovery, and stride-1 convolutions in both encoder and decoder maintain temporal continuity without aliasing artifacts. We have
@@ III-B MoSA-VQ: Motion Scale-Adaptive Residual VQ-VAE
Our full training objective combines four terms: motion reconstruction, codebook commitment, lightweight modulation regularization, and a novel cross-level decorrelation loss.
We adopt an reconstruction loss to ensure accurate motion recovery, and a standard commitment loss to encourage the encoder residuals to remain close to their quantized counterparts . A mild regularization is also applied to the per-layer scaling and bias parameters , following prior work on stable quantization with learnable modulation.
Additionally, we introduce a cross-level decorrelation loss, which explicitly penalizes statistical redundancy between the quantized vector and the subsequent residual :
This loss is motivated by the principle that, in an ideal residual quantization hierarchy, each level should encode novel information orthogonal to what has already been captured. A nonzero covariance between and indicates overlapping representation subspaces, leading to inefficient coding and excessive reliance on deeper quantizers. By minimizing their batch-wise covariance, we enforce an approximate statistical orthogonality between levels, which encourages disentangled information flow and hierarchical efficiency.
From an information-theoretic perspective, this decorrelation loss can be viewed as promoting minimal mutual information between consecutive encoding layers, thus maximizing the effective bit utilization of each quantized token. Empirically, we find this regularization yields more compact latent codes and improves reconstruction quality under constrained codebook budgets. The final loss function is:
Here, , and are weight factors.
III-C RQHC-Transformer: Residual Quantized Hierarchical Causal Transformer
To fully leverage the high-quality hierarchical representations produced by MoSA-VQ, we propose the Residual Quantized Hierarchical Causal Transformer (RQHC-Transformer), an autoregressive decoder structurally aligned with the residual quantization process. In contrast to single-resolution or per-layer decoding approaches, RQHC jointly models all quantization levels in a residual-aware and level-aware manner. Moreover, we adopt a causal Transformer design, which—compared to bidirectional transformers, LLM-style decoders, or iterative diffusion models—offers efficient one-pass inference, supports streamable generation, and naturally enables long-horizon synthesis. The coarse-to-fine decoding progression preserves structural semantics, mitigates token interference, and promotes stable, coherent motion generation with improved interpretability.
This hierarchical causal architecture progressively refines coarse motion structures, preserving semantic consistency across abstraction levels and reducing token-level interference. Crucially, its design is structurally aligned with the residual quantization process of MoSA-VQ, enabling stable information flow across layers. Together, this synergy allows MOGO to extend motion from any given frame with temporal continuity and semantic fidelity, laying the foundation for real-time, controllable infinite-length generation.
Hierarchical Causal Generation.
@@ III-C RQHC-Transformer: Residual Quantized Hierarchical Causal Transformer
where and are the query and key embeddings at position and , and is a learned relative position embedding, and and are learnable global bias vectors. The final output of the self-attention is:
@@
Training Objective.
To learn temporally coherent and hierarchically consistent token generation, we adopt a multi-layer autoregressive training objective defined over all quantization levels. The loss function is written as:
where denotes the total cross-entropy loss computed over all quantization levels. The training data consists of pairs , where is the set of discrete motion token sequences across quantization levels, and is the global conditioning signal derived from the input text and the residual context from coarser levels. Each sequence corresponds to the motion tokens at level , with being the number of time steps. The term denotes the causal context of , i.e., all tokens preceding the current position. The model , parameterized by weights , is trained to predict each token autoregressively based on its past tokens and the shared condition . This objective encourages consistent hierarchical structure and smooth temporal evolution across all levels.
III-D Infinite-Length Generation
By leveraging MoSA-VQ’s temporally discrete codes and RQHC’s autoregressive decoding, our framework naturally supports infinite-length motion generation from arbitrary time steps. During extension, the model conditions on previously generated tokens (excluding the original prompt embedding), while allowing the text prompt to be substituted dynamically—enabling seamless continuation under updated instructions.
Flexible Continuation with Prompt Switching.
At any generation step , given an existing motion token sequence at quantization level , we continue autoregressive generation from step onward using the RQHC-Transformer. To support runtime intervention, we allow prompt switching by replacing the original text condition with a new prompt , yielding an updated conditioning vector .
While the previous tokens are preserved, the updated prompt influences subsequent decoding. At each step , RQHC predicts the next token based on the causal context and the new condition:
This per-step decoding continues autoregressively and can be extended to arbitrary lengths, enabling real-time motion continuation from any point. Crucially, the updated prompt takes effect immediately from step , without requiring regeneration of prior tokens. Only after completing the desired generation (of any length) do we perform motion decoding to reconstruct full-body trajectories. This design enables flexible, prompt-controllable extension with temporal coherence and efficient computation.
Why This Works.
This flexible and prompt-adaptive generation capability stems from the close architectural synergy between MoSA-VQ and RQHC-Transformer. MoSA-VQ’s residual quantization provides a structured representation where coarse global motion is captured in early levels and finer dynamics are encoded hierarchically. This layered encoding ensures stability during long-horizon generation, as new content naturally extends existing motion without disrupting overall structure. In advanced, RQHC’s hierarchical causal structure enables prompt substitution at any point by conditioning future generation on preserved prior tokens, ensuring smooth semantic transitions. Together, these properties support real-time extension, runtime editing, and infinite-length motion generation under evolving conditions.
III-E Text Condition Alignment for Improved Motion Decoding
Real-world user prompts for motion generation often deviate from the structured annotations seen during training, varying in style, verbosity, and semantic complexity. Notably, many instructions contain abstract or composite motions that are difficult to execute when treated as a single condition, posing a fundamental challenge to robust generation.
To bridge this gap, we propose TCA, an inference-time strategy that leverages a pretrained large language model to enhance both interpretability and executability of free-form prompts—without modifying model parameters. TCA performs two key functions: (1) style normalization, aligning user input with the model’s linguistic prior, and (2) instruction decomposition, splitting complex intents into atomic motion steps.
Formally, given an arbitrary user prompt , TCA first produces a normalized instruction , then applies Chain-of-Thought prompting over to decompose it as:
where encodes few-shot exemplars and reasoning patterns.
Each denotes a semantically atomic sub-instruction, enabling the motion model to synthesize sequences in a stepwise or joint manner with improved controllability, precision, and temporal coherence.
By externally structuring complex instructions into executable units, TCA enhances generation quality across both seen and unseen prompts—without requiring any fine-tuning or additional supervision.